Westerville Company reported the following results from last year’s operations:

Sales $ 1,000,000
Variable expenses 300,000
Contribution margin 700,000
Fixed expenses 500,000
Net operating income $ 200,000
Average operating assets $ 625,000


At the beginning of this year, the company has a $120,000 investment opportunity with the following cost and revenue characteristics:

Sales $ 200,000
Contribution margin ratio 60 % of sales
Fixed expenses $ 90,000

The company’s minimum required rate of return is 15%.

Foundational 10-10

10-a. If Westerville’s chief executive officer will earn a bonus only if her ROI from this year exceeds her ROI from last year, would she pursue the investment opportunity?

  • Yes

  • No

10-b. Would the owners of the company want her to pursue the investment opportunity?

  • Yes

  • No

11. What is last year’s residual income?

12. If Westerville’s chief executive officer will earn a bonus only if her residual income from this year exceeds her residual income from last year, would she pursue the investment opportunity?

  • Yes

  • No

15-a. Assume that the contribution margin ratio of the investment opportunity was 50% instead of 60%. If Westerville’s chief executive officer will earn a bonus only if her residual income from this year exceeds her residual income from last year, would she pursue the investment opportunity?

  • Yes

  • No

15-b. Would the owners of the company want her to pursue the investment opportunity?

  • Yes

  • No
Answer #1

10 (a)

ROI from last year = 200,000 / 625,000 * 100 = 32%

ROI from new investment of $120,000 = [(200,000 * 60%) - 90,000] / 120,000 * 100= 25%

Thus. the new investment will lead to a Dilution in ROI if accepted. Hence, westervilles's CEO should NOT pursue the investment opportunity.

10 (b)

Company's minimum required rate of return = 15%

Rate of return promised by the new investment = 25% (calculated above)

As the actual rate of return expected exceeds the minimum required rate of return, the owners of the company would want to pursue the investment opportunity

11.

Residual Income refers to the excess income generated more than the minimum required rate of return.

Residual income = Net operating income - (minimum required rate of return * operating assets)

= 200,000 - (15% * 625,000)

= $ 106,250

12.

Net operating income from new investment of $ 120,000 = [(200,000 * 60%) - 90,000 = $30,000

Residual income if investment opportunity is accepted

= (200,000 + 30,000) - (15% * [625,000 + 120,000])

= 230,000 - (15% * 745,000)

= 118,250

As the residual income from this year ($118250) exceeds the residual income from last year ($106250), the CEO should pursue the investment opportunity

15. (a)

Net operating income from new investment if contribution margin is 50% = [(200,000 * 50%) - 90,000 = $10,000

Residual income if investment opportunity is accepted

= (200,000 + 10,000) - (15% * [625,000 + 120,000])

= 210,000 - (15% * 745,000)

= 98,250

As the residual income from this year ($98250) is less than the residual income from last year ($106250), the CEO should NOT pursue the investment opportunity

15. (b)

ROI from new investment if contribution margin is 50%

= [(200,000 * 50%) - 90,000] / 120,000 * 100= 8.33%

As the actual rate of return expected (8.33%) is less than the minimum required rate of return (15%), the owners of the company would NOT want to pursue the investment opportunity